P 13-17 (similar to) Question Help Growth Companys current share price is $19.95 and it is expected to pay a $1.25 dividend per share next year. After that, the firms dividends are expected to grow at a rate of 3.5% per year a. What is an estimate of Growth Companys cost of equity? b. Growth Company also has preferred stock outstanding that pays a $2.25 per share fixed dividend. If this stock is currently priced at $27.90, what is Growth Companys cost of preferred stock? C. Growth Company has existing debt issued three years ago with a coupon rate of 5.5%. The firm just issued new debt at par with a coupon rate of 6.2%. What is Growth Companys cost of debt? d. Growth Company has 4.9 million common shares outstanding and 1.3 million preferred shares outstanding, and its equity has a total book value of $50.0 million. Its liabilities have a market value of $19.7 million. If Growth Companys common and preferred shares are priced as in parts (a) and (b), what is the market value of Growth Companys assets? e. Growth Company faces a 38% tax rate. Given the information in parts (a) through (d), and your answers to those problems, what is Growth Companys WACC? a. What is an estimate of Growth Companys cost of equity? The required return (cost of capital) of levered equity is %. (Round to two decimal places.)
Answer #1

Di P(Stock Price1 T-g

where D1 = next year's dividend

r = cost of equity

g = dividend growth rate

=> r = \frac{D_{1}}{P}+g

r = \frac{1.25}{19.95}+0.035

r = 0.09765 = 9.765%

b)

Cost of Preferred Stock = Annual Preferred Dividend / Price of preferred stock

= $2.25 / $27.9

= 8.064%

c) Growth company's cost of debt = 6.2%, which is the coupon payment rate of the latest bond issue

d) Market Value of Assets = Market Value of Equity + Market Value of Liabilities

Market Value of Equity = no. of common shares * common stock price + no. of preferred shares * price of preferred shares

= 4.9 million * $19.95 + 1.3 million * $27.9

=$97.755 million + $36.27 million

=$134.025 million

and market value of liabilies is given as $19.7 million

Market Value of Assets = $ 134.025 million + $19.7 million = $153.725 million

e)

WACC = cost of equity*proportion of common equity + cost of preference equity * proportion of preference equity + cost of debt (1-t) * proportion of debt

total value of company = Market Value of Assets = $153.725 million

proportion of common equity = $97.755 million / $153.725 million = 0.6359

proportion of preference equity = $36.27 million / $153.725 million = 0.2359

proportion of debt = 1 - 0.2359 - 0.6359 = 0.1282

WACC = 0.6359 * 9.765% + 0.2359 * 8.064% + 0.1282 * (1-0.38)*6.02%

= 6.2095% + 1.9022% + 0.4784%

=8.5901%
Answer #2

DA SILVA Plc’s shares are not traded in any recognized market. Its sole activity is saloon and car hire. It is financed by a combination of 2 million $0.5 ordinary shares and a $1.5 bank loan. Very recently, Mavi’s Plc, a national car hire group offered a total of $5.5 million to acquire equity of DA SILVA Plc. The bid failed because majority of shareholders rejected since they wished to retain control of business, despite believing the offer to represent a fair price for the shares. The bank loan is at a floating rate of 10% p.a. and is secured on various fixed assets. The value of the bank loan is considered to be very close to its nominal value.

DA SILVA Plc’s current capital structure (by market value) represents what has been and is intended to continue to be its target capital structure.

DA SILVA Plc’s management is in the process of assessing a major investment, to be financed from retained earnings, in some new deposits, similar to the business’s existing ones. An appropriate cost of capital is required to this purpose. The dividend growth model has been proposed as a suitable basis for the estimation of equity.

Recent annual dividends per share have been:

Year            $

1                 0.0800

2                 0.0900

3                 0.1050

4                 0.1125

5                 0.1250

6                 0.1350

7                 0.1450

8                 0.1550

Corporation tax rate is expected to be 33% p.a. for the foreseeable future.

Required:

Estimate:

a)      Growth rate, g

b)     Cost of equity, Ke

c)                    

 


answered by: ABASS